One number is 6 more than twice another. Their product is 6 more than twice their sum. Find the numbers x and y
Answer by ewatrrr(24785)

 
Hi,
One number is 6 more than twice another
Let x and ( 2x+6) represent the numbers
Question states***
x(2x+6) = 2[x + (2x+6)] + 6
Solving for x
2x^2 + 6x = 6x + 18
2x^2 = 18
x^2 = 9
x = ± 3
Numbers are 3 and 12 OR -3 and 0
CHECKING our Answer***
3*12 = 2*15 + 6 = 36
-3*0 = 0 = 2*(-3) + 6